What are the dimensions of the largest rectangular field which can be enclosed by 1200 cm of fencing?
Can you please help me out and also show the work ? Thanks so much in advance:)

Thanks so much in advance:)

Answer by TimothyLamb(4379)

---
2L + 2w = 1200
2L = 1200 - 2w
L = 600 - w
a = Lw
a = (600 - w)w
a = -ww + 600w
---
the above quadratic equation is in standard form, with a=-1, b=600, and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 600 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic vertex is a maximum at: ( w= 300, a= 90000 )
---
answer:
a maximum area of 90,000 sq.cm is enclosed by:
L = 300 cm
w = 300 cm
---
the zero root doesn't fit the problem statement, so use the other root:
---
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php